# Towards a RAG based Summarization for Electron Ion Collider

### v1

@@ 1 Background
The EIC is the next generation large-scale scientific project that is anticipated to become operational within the coming decade [1]. The EIC User group is a fast-evolving community consisting of more than 1,400 physicists from over 38 countries around the world. Book keeping has been a significant challenge in such large collaborative experiments; specifically, in the case of large-scale physics experiments, one faces significant challenges in coordinating information curation between numerous working groups within the collaboration.
@@
Fine-tuning an LLM involves refining its abilities and performance in specific tasks or domains by training it further in domain-specific datasets after pretraining, improving effectiveness without retraining the entire model [12, 10]. However, conventional fine-tuning requires substantial computational power and time, posing challenges, especially for extensive models such as GPT-3 (175-B parameters)[3] and Meta LLaMA2 (13-B parameters) [13]. Despite the advent of Low-Rank Adaptation (LoRA) which enables efficient fine-tuning, even on consumer-grade GPUs, fine-tuning remains computationally intensive. Integrating strategies like in-context learning and chain-of-thought techniques [4] with a live knowledge repository forms the concept of Retrieval Augmented Generation (RAG), reducing hallucinations and anchoring LLMs to reality.
@@ 2 Retrieval Augmented Generation pipeline
Figure 1: A Naive RAG Agent involves the user inputting a prompt to the agent. The agent leverages an external knowledge base to gather more details related to the query. The information is enhanced using a predefined response template, which is inputted into a Frozen LLM such as OpenAI’s GPT, Anthropic’s Claude or Meta’s LLaMA2 for summarization of the result. The RAG Agent operates within a LLM engineering platform such as LangChain, LangFuse or LaamaIndex frameworks.
@@ 3 Creation of Knowledge base
The arXiv dataset:
We built a selected collection of papers pertaining to the Electron Ion Collider from the arXiv database. These articles specifically cover the period of time after 2021. This approach was taken to ensure that the language model (ChatGPT 3.5) did not possess prior knowledge of the content contained within these articles, as its training data only extend up to September 2021 [11]. Consequently, any new information produced after this date is unlikely to be implicitly stored in the model weights. The complete list of arXiv sources utilized for this study can be accessed at ARXIV_SOURCES.info.
@@
Other local database such as
LanceDB and ChromaDB was explored
@@
Table 1: Table presenting the specifics of the ingestion pipeline for the created RAG Agent. Various methodological investigations have been carried out and are currently in progress, influencing the RAG Agent’s performance. PineCone VectorDatabase was selected for its simplicity in deployment and utilization throughout the ingestion and inference stages. More details on the ingestion process can be found in the developed web application https://rags4eic-ai4eic.streamlit.app/.
It is crucial to emphasize that each entry in Table. 1 has an impact on the precision of vector retrieval. The cosine similarity metric is utilized to assess the degree of similarity between two sets of semantic information. The efficacy of this metric is closely related to the size of the data segments, with a reduction in the effectiveness of cosine similarity as the segment size increases. On the contrary, if the segment sizes are too small, there is the possibility of losing the semantic relevance of the data segments. Optimizing parameters is crucial for enhancing vector retrieval efficiency, with performance measured by retrieval latency and similarity metrics, involving inherent trade-offs and objectives. As this pilot study aims to demonstrate RAG’s feasibility for EIC, detailed optimization is deferred to future work.
@@ 4 Inference
Figure 3: When a user submits a question and selects both the metric and search configuration for vector searching, the query first passes through a decision chain. The LLM evaluates whether to consult the knowledge base for the answer. If more information is needed, it searches the vector database for additional context and sources. This information is processed through a fine-tuned template to craft a response with citations. The LLM then ensures syntax accuracy, delivering a GitHub markdown formatted answer. ChatGPT-3.5 functions as the LLM in this workflow.
@@
  • Instruction tuned prompts: To enhance the alignment of LLM for particular tasks, contextual learning techniques are used. Prompts that contain sample responses have been shown to improve summary performance. The prompt templates utilized for this study are outlined in the GitHub repository https://github.com/ai4eic/EIC-RAG-Project/tree/main/Templates.
@@
5 Evaluation of the developed RAG
To assess the effectiveness of the RAG Agent, it is necessary to create a dataset comprising a series of questions along with their optimal answers and sources. This process typically requires a significant amount of human resources due to the need for expertise in the field. The dataset selected for this research encompasses a variety of disciplines, ranging from hep.ph to nucl-ex to ph-acc [^2] These are the arXiv categorical codes in physics. This inherently requires domain expertise to create reliable QA datasets. However, to alleviate the requirement for domain-specific knowledge, an AI-driven approach was adopted to produce a high quality dataset. The created dataset consists of targeted questions with a complete response along with the source information.
@@ 5 Evaluation of the developed RAG
The primary goal of developing this data set is to utilize an advanced language model such as ChatGPT4.0 to generate a synthetic data set for evaluation purposes. Each question in the dataset is associated with a clearly defined set of "claims". Each question is uniquely characterized by the number of claims it addresses, and the corresponding answers to these claims are produced, culminating in a comprehensive response. In order to better facilitate this process, a routine within the web application was created where “annotators" can generate new Question and Answer pairs and add to datasets. The dataset creation process involves the following steps:
@@
    The “annotator” chooses an arXiv paper (with an option for a random, unexplored selection), the total questions to generate, and the claims per question. ChatGPT-4.0 then processes the paper’s contents using a template to produce formatted Question and Answer pairs.
@@
    Each QA pair has a question with ‘N’ claims and a detailed json object which has detailed information about the answers. The json object contains the number of claims in the questions, the individual claims, ideal response to each of the individual claims, and a complete response involving all the claims. Hence, this method of data generation provides a rich and high quality dataset for evaluation. This approach also guarantees that even someone new to the field can generate a dataset with the help of LLM.
@@
Each of the questions in the data set has a defined number of claims ($M$). Each of the questions has the source of the article, individual responses to the claims, and a complete ideal response. The performance of the RAG Agent, along with a brief description of the metric, is shown in Table 2.444The RAG Agent output response is captured as a PDF document and can be found here.
@@
The RAGAs framework was created to evaluate the effectiveness of QA chatbots [5]. This evaluation framework offers a comprehensive assessment of the performance of the RAG pipeline. A significant feature of the RAGA framework is its capability to adapt to the use of LLMs for evaluation without the need for reference annotations by human experts. This strategy not only simplifies the evaluation process but also improves scalability by reducing the dependence on manual annotations. However, RAGAs also have potential drawbacks, particularly in terms of their subjective nature in the evaluation. This subjectivity could lead to biases and inconsistencies in the evaluation process if not closely monitored. Therefore, a controlled application of RAGAs is crucial during the evaluation phase. The performance of the RAG Agent on the RAGAs score is summarized in Table 3555LLM used by RAGAs to score is ChatGPT4.0..
@@
Key observations and conclusions on the performance of the RAG Agent are summarized in the following:
@@ 6 Conclusions
A RAG Agent has been successfully developed as a pilot study case for the upcoming EIC. The RAG Agent has demonstrated its capability to condense data sets and information and provide trustworthy references with very few instances of hallucinations, enhancing the understanding and effective information retrieval for the EIC community. A complete web application has been developed from dataset creation to inference. Detailed tracing is currently necessary for better tracking of these chains using LangSmith. The study has successfully demonstrated strong motivation to continue developing modular RAG methods for effective information retrieval for EIC. The use of LLM for creating artificial data has been studied; however, there is a requirement to create datasets by experts in the field to create top-notch reference data for evaluating the RAG pipelines. By improving an LLM with these datasets, it is possible to improve the techniques of LLM-supported synthetic data generation, providing a scalable method to generate extensive benchmark datasets. The subsequent phase in this area involves engaging diverse domain specialists to begin using the web tool and efficiently gather high-quality data from their use. The software infrastructure for these efforts is nearing its mature phase. Evaluation pipelines such as RAGAs have been utilized to assess the performance of the RAG Agent without the need for references. The evaluators have not undergone parameter optimization to prevent any impact on the scores obtained. Despite the inherent challenge of reproducibility, efforts are being made to address this issue by effectively tracing the logic chains with tools such as LangSmith in order to reproduce the results on references. Multiple systematic investigations are needed to improve the comprehension of how the performance of the RAG system is influenced by factors such as ingestion, retrieval chain, LLM (e.g., ChatGPT3.5, ChatGPT4.0, Claude, LLaMA2), and potentially other hyperparameters.

### v3

# Towards a RAG-based Summarization for the Electron Ion Collider

@@ 1 Background
The EIC is the next generation large-scale scientific project that is anticipated to become operational within the next decade [1]. The EIC User group is a fast-evolving community consisting of more than 1,400 physicists from over 38 countries around the world. Book keeping has been a significant challenge in such large collaborative experiments; specifically, in the case of large-scale physics experiments, one faces significant challenges in coordinating information curation between numerous working groups within the collaboration.
@@
Fine-tuning an LLM involves refining its abilities and performance in specific tasks or domains by training it further in domain-specific datasets after pretraining, improving effectiveness without retraining the entire model [12, 10]. However, conventional fine-tuning requires substantial computational power and time, posing challenges, especially for extensive models such as GPT-3 (175-B parameters)[3] and Meta LLaMA2 (13-B parameters) [13]. Despite the advent of Low-Rank Adaptation (LoRA) which enables efficient fine-tuning, even on consumer-grade GPUs, fine-tuning remains computationally intensive. Integrating strategies like in-context learning and chain-of-thought techniques [4] with a live knowledge repository forms the concept of Retrieval Augmented Generation (RAG), reducing hallucinations and anchoring the LLMs to reality
@@ 2 Retrieval Augmented Generation pipeline
Figure 1: A Naive RAG Agent involves the user inputting a prompt to the agent. The agent leverages an external knowledge base to gather more details related to the query. The information is enhanced using a predefined response template, which is fed in as an input to a Frozen LLM such as OpenAI’s GPT, Anthropic’s Claude or Meta’s LLaMA2 for summarization of the result. The RAG Agent operates within a LLM engineering platform such as LangChain, LangFuse or LlamaIndex frameworks.
@@ 3 Creation of Knowledge base
EIC arXiv dataset:
We built a selected collection of papers pertaining to the Electron Ion Collider from the arXiv database. These articles specifically cover the period of time after 2021. This approach was taken to ensure that the language model (GPT 3.5) did not possess prior knowledge of the content contained within these articles, as its training data only extend up to September 2021 [11]. Consequently, any new information produced after this date is unlikely to be implicitly stored in the model weights. The complete list of arXiv sources utilized for this study can be accessed at ARXIV_SOURCES.info.
@@
Other local databases such as LanceDB and ChromaDB were explored
@@
Table 1: Table presenting the specifics of the ingestion pipeline for the created RAG Agent. Various methodological investigations have been carried out and are currently in progress, influencing the RAG Agent’s performance. PineCone VectorDatabase was selected for its simplicity in deployment and utilization throughout the ingestion and inference stages. More details on the ingestion process can be found in the tagged version of the source code on GitHub.
It is crucial to emphasize that each entry in Table. 1 has an impact on the precision of vector retrieval. The cosine similarity metric is utilized to assess the degree of similarity between two sets of semantic information. The efficacy of this metric is closely related to the size of the data segments, with a reduction in the effectiveness of cosine similarity as the segment size increases. On the contrary, if the segment sizes are too small, there is the possibility of losing the semantic relevance of the data segments. Optimizing parameters is crucial for enhancing vector retrieval efficiency, with performance measured by retrieval latency and similarity metrics, involving inherent trade-offs and objectives. As this pilot study aims to demonstrate the feasibility of RAG for the EIC, detailed optimization is deferred to future work.
@@ 4 Inference
Figure 3: When a user submits a question and selects both the metric and search configuration for vector searching, the query first passes through a decision chain. The LLM evaluates whether to consult the knowledge base for the answer. If more information is needed, it searches the vector database for additional context and sources. This information is processed through a fine-tuned template to craft a response with citations. The LLM then ensures syntax accuracy, delivering a GitHub markdown formatted answer. ChatGPT was used for this workflow with GPT-3.5 (gpt-3.5-turbo-1106) as the LLM running on the backend.
@@
  • Instruction tuned prompts: To enhance the alignment of LLM for particular tasks, contextual learning techniques are used. Prompts that contain sample responses have been shown to improve summary performance. The prompt templates utilized for this study are outlined in the tagged source code on the GitHub repository.
@@
5 Evaluation of the EIC-RAG
To assess the effectiveness of the RAG Agent, it is necessary to create a dataset comprised of series of questions along with their optimal answers and sources. This process typically requires a significant amount of human resources due to the need for expertise in the field. The dataset selected for this research encompasses a variety of disciplines, ranging from hep.ph to nucl-ex to ph-acc [^2] These are the arXiv categorical codes in physics. This inherently requires domain expertise to create reliable QA datasets. However, to alleviate the requirement for domain-specific knowledge, an AI-driven approach was adopted to produce a high quality dataset. The created dataset consists of targeted questions with a complete response along with the source information.
@@ 5 Evaluation of the EIC-RAG
The primary goal of developing this data set is to utilize an advanced language model such as GPT-4.0 to generate a synthetic data set for evaluation purposes. Each question in the dataset is associated with a clearly defined set of "claims". Each question is uniquely characterized by the number of claims it addresses, and the corresponding answers to these claims are produced, culminating in a comprehensive response. In order to better facilitate this process, a routine within the web application was created where “annotators" can generate new Question and Answer pairs and add to datasets. The dataset creation process involves the following steps:
@@
    The “annotator” chooses an arXiv paper (with an option for a random, unexplored selection), the total questions to generate, and the claims per question. GPT-4.0 then processes the paper’s contents using a template to produce formatted Question and Answer pairs.
@@
    Each QA pair has a question with ‘N’ claims and a detailed json object which has detailed information about the answers. The json object contains the number of claims in the questions, the individual claims, ideal response to each of the individual claims, and a complete response involving all the claims. Hence, this method of data generation provides a rich and high quality dataset for evaluation. This approach also guarantees that even someone new to the field can generate a dataset with the help of LLMs.
@@
Each of the questions in the data set has a defined number of claims ($M$). Each of the questions has the source of the article, individual responses to the claims, and a complete ideal response. The Retrieval Augmented Generation Assessment, along with a brief description of the metric, is shown in Table 2.444The RAG Agent output response is captured as a PDF document and can be found here.
@@
The RAGAs framework was created to evaluate the effectiveness of QA chatbots [5]. This evaluation framework offers a comprehensive assessment of the performance of the RAG pipeline. A significant feature of the RAGA framework is its capability to adapt to the use of LLMs for evaluation without the need for reference annotations by human experts. This strategy not only simplifies the evaluation process but also improves scalability by reducing the dependence on manual annotations. However, RAGAs also have potential drawbacks, particularly in terms of their subjective nature in the evaluation. This subjectivity could lead to biases and inconsistencies in the evaluation process if not closely monitored. Therefore, a controlled application of RAGAs is crucial during the evaluation phase. The performance of the RAG Agent on the RAGAs score is summarized in Table 3555LLM used by RAGAs to score is GPT4.0..
@@
The principal observations and corresponding conclusions regarding the performance of the RAG Agent are delineated below. Additionally, the subsequent steps to enhance the efficacy of EIC-RAG are enumerated.:
@@ 6 Conclusions
A RAG Agent has been successfully developed as a pilot study case for the upcoming EIC. The RAG Agent has demonstrated its capability to condense data sets and information and provide trustworthy references with very few instances of hallucinations, enhancing the understanding and effective information retrieval for the EIC community. A complete web application has been developed from dataset creation to inference. Detailed tracing is currently necessary for better tracking of these chains using LangSmith. The study has successfully demonstrated strong motivation to continue developing modular RAG methods for effective information retrieval for EIC. The use of LLM for creating artificial data has been studied; however, there is a requirement to create datasets by experts in the field to create top-notch reference data for evaluating the RAG pipelines. By improving an LLM with these datasets, it is possible to improve the techniques of LLM-supported synthetic data generation, providing a scalable method to generate extensive benchmark datasets. The subsequent phase in this area involves engaging diverse domain specialists to begin using the web tool and efficiently gather high-quality data from their use. The software infrastructure for these efforts is nearing its mature phase. Evaluation pipelines such as RAGAs have been utilized to assess the performance of the RAG Agent without the need for references. The evaluators have not undergone parameter optimization to prevent any impact on the scores obtained. Despite the inherent challenge of reproducibility, efforts are being made to address this issue by effectively tracing the logic chains with tools such as LangSmith in order to reproduce the results on references. Multiple systematic investigations are needed to improve the comprehension of how the performance of the RAG system is influenced by factors such as ingestion, retrieval chain, LLM (e.g., GPT3.5, GPT4.0, Claude, LLaMA2), and potentially other hyperparameters.
@@
Acknowledgements
K.S. is supported by the Office of Nuclear Physics of the U.S. Department of Energy under Grant Contract No. DE-SC0024625